The analytical framework used to evaluate transactions is reproduced below:

Cash

+

Non-Cash

Assets

=

Liabilities

+

Contributed

Capital

+

Accumulated Other

Comprehensive

Income

+

Retained

Earnings

Using this analytical framework, indicate the effect of each of the following transactions for Staples Corporation:

1.

Staples recorded cash sales of $25,000. The merchandise had cost $19,000 to manufacture.

2.

Staples purchased $8,500 of raw material inventory on account.

3.

The company paid $2,500 for property insurance for the next 12 months.

4.

Staples paid its employees $5,000 for the month.

5.

The company purchased $1,000 of supplies on account.

6.

Staples issued $25,000 of long-term debt.

7.

The company used $10,000 of excess cash to purchase marketable securities.

8.

Staples purchased a machine for $16,000 using $8,000 cash with the balance on account.

9.

Staples paid $2,500 for interest expense on the long-term debt.

10.

At the end of the year the marketable securities that Staples purchased in transaction 7 were now worth $14,500.

11.

Depreciation for the period was $1,500.

12.

Staples examined the equipment and determined that its fair value was $10,000.

ANS:

Cash

+

Non-Cash

Assets

=

Liabilities

+

Contributed

Capital

+

Accumulated Other

Comprehensive

Income

+

Retained

Earnings
Answer #1
Cash + Non-Cash Assets = Liabilities + Contributed Capital + Accumulated Other Comprehensive Income + Retained Earnings
1) $25,000 + = + + $25,000 +
+ ($19,000) = + + ($19,000) +
2) + $8,500 = $8,500 + + +
3) ($2,500) + $2,500 = + + +
4) ($5,000) + = + + ($5,000) +
5) + $1,000 = $1,000 + + +
6) $25,000 + = $25,000 + + +
7) ($10,000) + $10,000 = + + +
8) ($8,000) + $16,000 = $8,000 + + +
9) ($2,500) + = + + ($2,500) +
10) + $4,500 = + + $4,500 +
11) + ($1,500) = + + ($1,500) +
12) + ($6,000) = + + ($6,000) +
$22,000 + $16,000 = $42,500 + $0 + ($4,500) + $0